# Beer Path Problems in Temporal Graphs

### v1

@@ 2.1 Preliminaries
A temporal path $P$ in a temporal graph $G$ is a sequence of vertices $P=\langle v_{1},v_{2},\dots,v_{\ell}\rangle$ where $(v_{i},v_{i+1},t_{i},\lambda_{i})\in E$ is the $i$-th temporal edge on $P$ for $1\leq i\leq\ell$, and $(t_{i}+\lambda_{i})\leq t_{i+1}$ for $1\leq i<\ell$. The quantity $(t_{\ell}+\lambda_{\ell})$ is known as the ending time of $P$, denoted by $end(P)$, while $t_{1}$ is the starting time of $P$, $start(P)$. The duration of $P$ is defined as $dura(P)=end(P)-start(P)$. Finally, the sum of the traversal time of each temporal edge of $P$ is known as the distance of the path, denoted as $dist(P)$.
@@
Let $P$ be a temporal $u-v$ path. We say that $P$ is dominated if there exists another temporal $u-v$ path $P^{\prime}$ such that $\text{start}(P^{\prime})\geq\text{start}(P)$ and $\text{end}(P^{\prime})\leq\text{end}(P)$, with at least one of the inequalities being strict. We say that $P$ is distance-wise dominated if there exists a temporal $u-v$ path $P^{\prime}$ such that $\text{dist}(P^{\prime})\leq\text{dist}(P)$ and $\text{end}(P^{\prime})\leq\text{end}(P)$, with at least one of the inequalities being strict.
@@
1:A temporal graph $G=(V,E)$, source vertex $x$, time interval $[t_{\alpha},t_{\omega}]$.
2:For $v\in V$, the list $L_{v}$ of non-dominated $(s,a)$ pairs (from $x$ to $v$) within $[t_{\alpha},t_{\omega}]$.
@@
1:A temporal graph $G=(V,E)$, source vertex $x$, time interval $[t_{\alpha},t_{\omega}]$.
2:For $v\in V$, the list $S_{v}$ of distance-wise non-dominated $(d,a)$ pairs (from $x$ to $v$) within $[t_{\alpha},t_{\omega}]$.
@@ 2.2 The temporal beer path problem
Given a temporal graph $G$, vertices $x,y\in V$ and a time interval $[t_{\alpha},t_{\omega}]$, we denote the set of temporal beer paths between two vertices $x$ and $y$ starting after $t_{\alpha}$ and ending before $t_{\omega}$ as $P_{B}(x,y,[t_{\alpha},t_{\omega}])=$$\{P=\langle x=v_{1},v_{2},\dots,v_{\ell}=y\rangle:P$ is a temporal path from $x$ to $y$ such that $start(P)\geq t_{\alpha},\ end(P)\leq t_{\omega},\ \exists\ (i\in[\ell],t\in T_{v_{i}}):(v_{i}\in B,\ t\in[t_{i}+\lambda_{i},t_{i+1}])\}$. Note that a beer path must traverse at least one beer vertex at a time in which it is active.
@@
Figure 1: Example in which the $x-y$ EABP/LDBP/FBP/SBP use $b_{1},\ldots,b_{4}$ respectively. Numbers in parentheses denote $(t,\lambda)$ of edges and numbers in brackets denote beer vertex active times. $x-b_{5}-y$ is not a valid temporal beer path because $4\notin T_{b_{5}}$. The $x-b_{2}$ edge with $(t,\lambda)=(18,9)$ is dominated and can be removed from the graph without affecting any objective.
@@ 3.1.1 An earliest-arrival beer path algorithm in edge stream representation
Given a temporal graph $G=(V,E)$ and initial times $init[v]\in\mathbb{R}^{+}\cup\{\infty\}$, $v\in V$, we say that a temporal path $x-y$ starting from $x$ no earlier than $init[x]$ is a multiple-source earliest-arrival path to $y$ if its arrival time at $y$ is minimum among all temporal paths $v-y$, $v\in V$, starting from $v$ no earlier than $init[v]$. We call this arrival time the multiple-source earliest-arrival path time of $y$.
@@
Algorithm 3 MSEAP$(G,[t_{\alpha},t_{\omega}],init)$
1:A temporal graph $G$ in its edge stream representation, a time window $[t_{\alpha},t_{\omega}]$ and initial times $init[v]$, such that $t_{\alpha}\leq init[v]\leq t_{\omega}$ or $init[v]=\infty$, for all $v\in V$.
2:The MSEAP times within $[t_{\alpha},t_{\omega}]$ to every vertex, assuming we can start at time $init[v]$ from each $v\in V$.
3:$\tau[v]\leftarrow init[v]$ for all $v\in V$
@@
For all nodes $x,y\in V$, if there exists some temporal $x-y$ path, then there exists some earliest-arrival $x-y$ path $\langle v_{0}=x,v_{1},v_{2},\ldots,v_{k}=y\rangle$ for which every prefix-subpath $\langle x,v_{1},v_{2},\ldots,v_{i}\rangle$ is an earliest-arrival $x-v_{i}$ path.
@@
In temporal graphs, the property described in Lemma 3.2 only holds for some $x-y$ earliest-arrival path, in contrast to classical graphs, where a similar property holds for all $x-y$ shortest paths.
@@
Algorithm 3 computes the MSEAP times to each vertex, within time window $[t_{\alpha},t_{\omega}]$. If time $\infty$ is returned for some vertex, then that vertex is unreachable from all $v\in V$ (starting no earlier than $init[v]$) within $[t_{\alpha},t_{\omega}]$.
@@
For each $v\in V$ let $\tau[v]$ be the time calculated by the algorithm and $\tau^{*}[v]$ be the respective MSEAP time (within $[t_{\alpha},t_{\omega}]$). Since $\tau[v]$ is initialized to $init[v]$, any update done by the algorithm calculates the arrival time of a valid temporal path, starting from some node $v$ at time $init[v]$. Thus, at any step of the algorithm it holds that $\tau[v]\geq\tau^{*}[v]$, $\forall v\in V$.
@@
From the above we infer that $\tau[v]=\infty$ can be returned for some $v\in V$ only if no MSEAP to $v$ exists. This implies that $v$ is unreachable from all $u\in V$ (starting no earlier than $init[u]$) within $[t_{\alpha},t_{\omega}]$.
@@
Algorithm 4 EABP$(G,x,[t_{\alpha},t_{\omega}])$
1:A temporal graph $G$, source vertex $x$, time window $[t_{\alpha},t_{\omega}]$, a set of beer vertices $B$ and a list $T_{i}$ of time units for each of them.
2:The EABP time from $x$ to every vertex $v\in V$ within $[t_{\alpha},t_{\omega}]$.
3:$init[x]\leftarrow t_{\alpha},\,init[v]\leftarrow\infty$ for all $v\in V\setminus\{x\}$
4:$\tau[v]\leftarrow\texttt{MSEAP($G,[t_{\alpha},t_{\omega}],init$)}$, for $v\in V$
5:$init^{\prime}[v]\leftarrow\infty$ for all $v\in V$
@@
7:     $beer\_time\leftarrow\min\{time\in T_{b}\mid time\geq\tau[b]\}$triangleright $\infty$ if it does not exist
@@
10:$\tau^{\prime}[v]\leftarrow\texttt{MSEAP($G,[t_{\alpha},t_{\omega}],init^{\prime}$)}$, for $v\in V$
@@
By Observation 3.1.1, the first call to Algorithm 3 correctly computes the earliest-arrival path times from $x$ to the beer vertices (within $[t_{\alpha},t_{\omega}]$). Thus, the binary search on $T_{b}$ computes for each $b\in B$ the earliest departure time from $b$ s.t. $b$ is used as the beer vertex of a beer path from $x$ to any other node (within $[t_{\alpha},t_{\omega}]$). Denote this time as $t[b]$, $b\in B$. Observe that, for any node $v$, the EABP time from $x$ to $v$ is the earliest time in which $v$ is reachable from some $b\in B$, starting from $b$ no earlier than $t[b]$. This coincides with the MSEAP time to $v$ for $init[b]=t[b]$, $b\in B$, and $init[u]=\infty$, $u\in V\setminus B$. Hence, the second call to Algorithm 3 correctly computes the EABP times from $x$ to all nodes (within $[t_{\alpha},t_{\omega}]$), by Lemma 3.5.
@@ 3.1.2 A faster earliest-arrival beer path algorithm with adjacency list representation
Algorithm 5 MSEAP_alternative$(G,[t_{\alpha},t_{\omega}],init)$
1:A temporal graph $G$ in adjacency list representation, a time window $[t_{\alpha},t_{\omega}]$ and initial times $init[v]$, such that $t_{\alpha}\leq init[v]\leq t_{\omega}$ or $init[v]=\infty$, for all $v\in V$.
2:The MSEAP times within $[t_{\alpha},t_{\omega}]$ to every vertex, assuming we can start at time $init[v]$ from each $v\in V$.
3:$\tau[v]\leftarrow init[v]$ for all $v\in V$
4:$PQ\leftarrow\emptyset$triangleright a min-priority queue
5:$PQ.enqueue(\tau[v],v)$ for all $v\in V$triangleright Add $v$ to priority queue with $\tau[v]$ priority
@@ 3.2 Algorithms for latest-departure beer path
Given a temporal graph $G=(V,E)$ and final times $fin[v]\in\mathbb{R}^{+}\cup\{-\infty\}$, $v\in V$, we say that a temporal $x-y$ path arriving at $y$ no later than $fin[y]$ is a multiple-target latest-departure path from $x$ if its starting time from $x$ is maximum among all temporal paths $x-v$, $v\in V$, arriving at $v$ no later than $fin[v]$. We call this starting time the multiple-target latest-departure path time of $x$.
The main algorithm (analogous to Algorithm 4) sets $fin[y]$ to $t_{\omega}$ and $fin[v]$ to $\infty$- ∞ for all other nodes. A MTLDP subroutine is used to calculate latest-departure times in the first phase and LDBP times in the second phase. A search in each list $T_{b}$ is done before the second phase, similar to Algorithm 4.
@@ 3.3 Algorithm for fastest beer path
If there is a temporal $x-y$ beer path within time interval $[t_{\alpha},t_{\omega}]$, then there is an $x-y$ FBP $\langle v_{0}=x,v_{1},\ldots,v_{j}=b,\ldots,v_{p}=y\rangle$ within $[t_{\alpha},t_{\omega}]$, traversing $b\in B$ at a time $t\in T_{b}$, such that its subpaths $\langle v_{0}=x,v_{1},\ldots,v_{j}=b\rangle$ and $\langle v_{j}=b,v_{j+1},\ldots,v_{p}=y\rangle$ are non-dominated paths.
@@
Algorithm 6 FBP$(G,x,[t_{\alpha},t_{\omega}])$
1:A temporal graph $G$ in edge stream representation, source vertex $x$, time interval $[t_{\alpha},t_{\omega}]$, a set of beer vertices $B$ and a list $T_{i}$ of time units for each of them.
2:The FBP duration from $x$ to every vertex $v\in V$ within $[t_{\alpha},t_{\omega}]$.
3:$L_{b}\leftarrow\texttt{Non-dom\textunderscore paths($G,x,[t_{\alpha},t_{\omega}]$)}$, for $b\in B$triangleright calculate non-dom paths with Alg 1
@@
6:         $a^{\prime}\leftarrow\min\{time\in T_{b}\mid time\geq a\}$triangleright $\infty$ if it does not exist
@@
10:              Remove $(s,a)$ from $L_{b}$triangleright impossible to get beer
@@
12:$L_{v}^{\prime}\leftarrow\texttt{Non-dom\textunderscore paths($G^{\prime},\epsilon,[t_{\alpha},t_{\omega}]$)}$, for $v\in V$
@@ 3.4 Algorithm for shortest beer path
Let $P$ be a temporal $u-v$ path. We say that $P$ is inverse-distance-wise dominated if there exists another temporal $u-v$ path $P^{\prime}$ such that $\text{dist}(P^{\prime})\leq\text{dist}(P)$ and $\text{start}(P^{\prime})\geq\text{start}(P)$, with at least one of the inequalities being strict.
@@
If there is a temporal $x-y$ beer path within time interval $[t_{\alpha},t_{\omega}]$, then there is a $x-y$ SBP $\langle v_{0}=x,v_{1},\ldots,v_{j}=b,\ldots,v_{p}=y\rangle$ within $[t_{\alpha},t_{\omega}]$, traversing $b\in B$ at a time $t\in T_{b}$, s.t. $\langle v_{0}=x,v_{1},\ldots,v_{j}=b\rangle$ is a distance-wise non-dominated path (within $[t_{\alpha},t_{\omega}]$) and $\langle v_{j}=b,v_{j+1},\ldots,v_{p}=y\rangle$ is an inverse-distance-wise non-dominated path (within $[t_{\alpha},t_{\omega}]$).
@@
1:A temporal graph $G=(V,E)$ in reverse edge stream representation, target vertex $y$, time interval $[t_{\alpha},t_{\omega}]$.
2:For each $v\in V$, the list $I_{v}$ of inverse-distance-wise non-dominated $(d,s)$ pairs (from $v$ to $y$) within $[t_{\alpha},t_{\omega}]$.
@@
Algorithm 7 computes all inverse-distance-wise non-dominated $(d,s)$ pairs from each $v\in V$ to $y$ within $[t_{\alpha},t_{\omega}]$, in time $\mathcal{O}(n+M\log(d_{out}))$.
@@
Algorithm 8 SBP$(G,x,y,[t_{\alpha},t_{\omega}])$
1:A temporal graph $G=(V,E)$ in edge stream representation (both regular and reverse), source vertex $x$, target vertex $y$, time interval $[t_{\alpha},t_{\omega}]$, a set of beer vertices $B$ and a list $T_{i}$ of time units for each of them.
2:The SBP distance from $x$ to $y$ within $[t_{\alpha},t_{\omega}]$.
3:$dist\leftarrow\infty$
4:$L_{b}\leftarrow\texttt{Dist\_Non-dom\_paths($G,x,[t_{\alpha},t_{\omega}]$)}$, for $b\in B$triangleright Alg 2 from $x$ to $b\in B$
5:$I_{b}\leftarrow\texttt{Inv\textunderscore dist\textunderscore Non-dom\textunderscore paths($G,y,[t_{\alpha},t_{\omega}]$)}$, for $b\in B$triangleright Alg 7 from $b\in B$ to $y$
@@
8:         $a^{\prime}\leftarrow\min\{time\in T_{b}\mid time\geq a\}$triangleright $\infty$ if it does not exist
@@
12:              Remove $(d,a)$ from $L_{b}$triangleright impossible to get beer
@@
14:         Find $(d^{\prime},s)$ in $I_{b}$ with minimum $s$ such that $s\geq a$triangleright find best fitting path

### v2

@@ 2.1 Preliminaries
A temporal path $P$ in a temporal graph $G$ is a sequence of vertices $P=\langle v_{1},v_{2},\dots,v_{\ell}\rangle$ where $(v_{i},v_{i+1},t_{i},\lambda_{i})\in E$ is the $i$-th temporal edge on $P$ for $1\leq i\leq\ell$, and $(t_{i}+\lambda_{i})\leq t_{i+1}$ for $1\leq i<\ell$. The quantity $(t_{\ell}+\lambda_{\ell})$ is known as the ending time of $P$, denoted by $end(P)$, while $t_{1}$ is the starting time of $P$, $start(P)$. The duration of $P$ is defined as dura(P)=end(P)start(P)dura(P)=end(P)-start(P). Finally, the sum of the traversal time of each temporal edge of $P$ is known as the distance of the path, denoted as $dist(P)$.
@@
Let $P$ be a temporal uvu-v path. We say that $P$ is dominated if there exists another temporal uvu-v path $P^{\prime}$ such that $\text{start}(P^{\prime})\geq\text{start}(P)$ and $\text{end}(P^{\prime})\leq\text{end}(P)$, with at least one of the inequalities being strict. We say that $P$ is distance-wise dominated if there exists a temporal uvu-v path $P^{\prime}$ such that $\text{dist}(P^{\prime})\leq\text{dist}(P)$ and $\text{end}(P^{\prime})\leq\text{end}(P)$, with at least one of the inequalities being strict.
@@
1:A temporal graph $G=(V,E)$, source vertex $x$, time interval $t_{\alpha},t_{\omega}]$.
2:For $v\in V$, the list $L_{v}$ of non-dominated $(s,a)$ pairs (from $x$ to $v$) within $t_{\alpha},t_{\omega}]$.
@@
1:A temporal graph $G=(V,E)$, source vertex $x$, time interval $t_{\alpha},t_{\omega}]$.
2:For $v\in V$, the list $S_{v}$ of distance-wise non-dominated $(d,a)$ pairs (from $x$ to $v$) within $t_{\alpha},t_{\omega}]$.
@@ 2.2 The temporal beer path problem
Given a temporal graph $G$, vertices $x,y\in V$ and a time interval $t_{\alpha},t_{\omega}]$, we denote the set of temporal beer paths between two vertices $x$ and $y$ starting after $t_{\alpha}$ and ending before $t_{\omega}$ as $P_{B}(x,y,[t_{\alpha},t_{\omega}])=$ $\{P=\langle x=v_{1},v_{2},\dots,v_{\ell}=y\rangle:P$ is a temporal path from $x$ to $y$ such that $start(P)\geq t_{\alpha},\ end(P)\leq t_{\omega},\ \exists\ (i\in[\ell],t\in T_{v_{i}}):(v_{i}\in B,\ t\in[t_{i}+\lambda_{i},t_{i+1}])\}$. Note that a beer path must traverse at least one beer vertex at a time in which it is active.
@@
Figure 1: Example in which the xyx-y EABP/LDBP/FBP/SBP use $b_{1},\ldots,b_{4}$ respectively. Numbers in parentheses denote $(t,\lambda)$ of edges and numbers in brackets denote beer vertex active times. $x-b_{5}-y$ is not a valid temporal beer path because $4\notin T_{b_{5}}$. The $x-b_{2}$ edge with $(t,\lambda)=(18,9)$ is dominated and can be removed from the graph without affecting any objective.
@@ 3.1.1 An earliest-arrival beer path algorithm in edge stream representation
Given a temporal graph $G=(V,E)$ and initial times $init[v]\in\mathbb{R}^{+}\cup\{\infty\}$, $v\in V$, we say that a temporal path xyx-y starting from $x$ no earlier than $init[x]$ is a multiple-source earliest-arrival path to $y$ if its arrival time at $y$ is minimum among all temporal paths vyv-y, $v\in V$, starting from $v$ no earlier than $init[v]$. We call this arrival time the multiple-source earliest-arrival path time of $y$.
@@
[…]
1:A temporal graph $G$ in its edge stream representation, a time window $t_{\alpha},t_{\omega}]$ and initial times $init[v]$, such that $t_{\alpha}\leq init[v]\leq t_{\omega}$ or $init[v]=\infty$, for all $v\in V$.
2:The MSEAP times within $t_{\alpha},t_{\omega}]$ to every vertex, assuming we can start at time $init[v]$ from each $v\in V$.
[…]
@@
For all nodes $x,y\in V$, if there exists some temporal xyx-y path, then there exists some earliest-arrival xyx-y path $\langle v_{0}=x,v_{1},v_{2},\ldots,v_{k}=y\rangle$ for which every prefix-subpath $\langle x,v_{1},v_{2},\ldots,v_{i}\rangle$ is an earliest-arrival $x-v_{i}$ path.
@@
In temporal graphs, the property described in Lemma 3.2 only holds for some xyx-y earliest-arrival path, in contrast to classical graphs, where a similar property holds for all xyx-y shortest paths.
@@
Algorithm 3 computes the MSEAP times to each vertex, within time window $t_{\alpha},t_{\omega}]$. If time $\infty$ is returned for some vertex, then that vertex is unreachable from all $v\in V$ (starting no earlier than $init[v]$) within $t_{\alpha},t_{\omega}]$.
@@
For each $v\in V$ let $\tau[v]$ be the time calculated by the algorithm and $\tau^{*}[v]$ be the respective MSEAP time (within $t_{\alpha},t_{\omega}]$). Since $\tau[v]$ is initialized to $init[v]$, any update done by the algorithm calculates the arrival time of a valid temporal path, starting from some node $v$ at time $init[v]$. Thus, at any step of the algorithm it holds that $\tau[v]\geq\tau^{*}[v]$, $\forall v\in V$.
@@
From the above we infer that $\tau[v]=\infty$ can be returned for some $v\in V$ only if no MSEAP to $v$ exists. This implies that $v$ is unreachable from all $u\in V$ (starting no earlier than $init[u]$) within $t_{\alpha},t_{\omega}]$.
@@
[…]
1:A temporal graph $G$, source vertex $x$, time window $t_{\alpha},t_{\omega}]$, a set of beer vertices $B$ and a list $T_{i}$ of time units for each of them.
2:The EABP time from $x$ to every vertex $v\in V$ within $t_{\alpha},t_{\omega}]$.
[…]
@@
7:  $beer\_time\leftarrow\min\{time\in T_{b}\mid time\geq\tau[b]\}$ $\triangleright$ $\infty$ if it does not exist
@@
[…]
@@
By Observation 3.1.1, the first call to Algorithm 3 correctly computes the earliest-arrival path times from $x$ to the beer vertices (within $t_{\alpha},t_{\omega}]$). Thus, the binary search on $T_{b}$ computes for each $b\in B$ the earliest departure time from $b$ s.t. $b$ is used as the beer vertex of a beer path from $x$ to any other node (within $t_{\alpha},t_{\omega}]$). Denote this time as $t[b]$, $b\in B$. Observe that, for any node $v$, the EABP time from $x$ to $v$ is the earliest time in which $v$ is reachable from some $b\in B$, starting from $b$ no earlier than $t[b]$. This coincides with the MSEAP time to $v$ for $init[b]=t[b]$, $b\in B$, and $init[u]=\infty$, $u\in V\setminus B$. Hence, the second call to Algorithm 3 correctly computes the EABP times from $x$ to all nodes (within $t_{\alpha},t_{\omega}]$), by Lemma 3.5.
@@ 3.1.2 A faster earliest-arrival beer path algorithm with adjacency list representation
[…]
1:A temporal graph $G$ in adjacency list representation, a time window $t_{\alpha},t_{\omega}]$ and initial times $init[v]$, such that $t_{\alpha}\leq init[v]\leq t_{\omega}$ or $init[v]=\infty$, for all $v\in V$.
2:The MSEAP times within $t_{\alpha},t_{\omega}]$ to every vertex, assuming we can start at time $init[v]$ from each $v\in V$.
[…]
@@ 3.2 Algorithms for latest-departure beer path
Given a temporal graph $G=(V,E)$ and final times $fin[v]\in\mathbb{R}^{+}\cup\{-\infty\}$, $v\in V$, we say that a temporal xyx-y path arriving at $y$ no later than $fin[y]$ is a multiple-target latest-departure path from $x$ if its starting time from $x$ is maximum among all temporal paths xvx-v, $v\in V$, arriving at $v$ no later than $fin[v]$. We call this starting time the multiple-target latest-departure path time of $x$.
The main algorithm (analogous to Algorithm 4) sets $fin[y]$ to $t_{\omega}$ and $fin[v]$ to -$\infty$ for all other nodes. A MTLDP subroutine is used to calculate latest-departure times in the first phase and LDBP times in the second phase. A search in each list $T_{b}$ is done before the second phase, similar to Algorithm 4.
@@ 3.3 Algorithm for fastest beer path
If there is a temporal xyx-y beer path within time interval $t_{\alpha},t_{\omega}]$, then there is an xyx-y FBP $\langle v_{0}=x,v_{1},\ldots,v_{j}=b,\ldots,v_{p}=y\rangle$ within $t_{\alpha},t_{\omega}]$, traversing $b\in B$ at a time $t\in T_{b}$, such that its subpaths $\langle v_{0}=x,v_{1},\ldots,v_{j}=b\rangle$ and $\langle v_{j}=b,v_{j+1},\ldots,v_{p}=y\rangle$ are non-dominated paths.
@@
[…]
1:A temporal graph $G$ in edge stream representation, source vertex $x$, time interval $t_{\alpha},t_{\omega}]$, a set of beer vertices $B$ and a list $T_{i}$ of time units for each of them.
2:The FBP duration from $x$ to every vertex $v\in V$ within $t_{\alpha},t_{\omega}]$.
[…]
@@
6:   $a^{\prime}\leftarrow\min\{time\in T_{b}\mid time\geq a\}$ $\triangleright$ $\infty$ if it does not exist
@@
10:     Remove $(s,a)$ from $L_{b}$ $\triangleright$ impossible to get beer
@@
[…]
@@ 3.4 Algorithm for shortest beer path
Let $P$ be a temporal uvu-v path. We say that $P$ is inverse-distance-wise dominated if there exists another temporal uvu-v path $P^{\prime}$ such that $\text{dist}(P^{\prime})\leq\text{dist}(P)$ and $\text{start}(P^{\prime})\geq\text{start}(P)$, with at least one of the inequalities being strict.
@@
If there is a temporal xyx-y beer path within time interval $t_{\alpha},t_{\omega}]$, then there is a xyx-y SBP $\langle v_{0}=x,v_{1},\ldots,v_{j}=b,\ldots,v_{p}=y\rangle$ within $t_{\alpha},t_{\omega}]$, traversing $b\in B$ at a time $t\in T_{b}$, s.t. $\langle v_{0}=x,v_{1},\ldots,v_{j}=b\rangle$ is a distance-wise non-dominated path (within $t_{\alpha},t_{\omega}]$) and $\langle v_{j}=b,v_{j+1},\ldots,v_{p}=y\rangle$ is an inverse-distance-wise non-dominated path (within $t_{\alpha},t_{\omega}]$).
@@
1:A temporal graph $G=(V,E)$ in reverse edge stream representation, target vertex $y$, time interval $t_{\alpha},t_{\omega}]$.
2:For each $v\in V$, the list $I_{v}$ of inverse-distance-wise non-dominated $(d,s)$ pairs (from $v$ to $y$) within $t_{\alpha},t_{\omega}]$.
@@
Algorithm 7 computes all inverse-distance-wise non-dominated $(d,s)$ pairs from each $v\in V$ to $y$ within $t_{\alpha},t_{\omega}]$, in time $\mathcal{O}(n+M\log(d_{out}))$.
@@
[…]
1:A temporal graph $G=(V,E)$ in edge stream representation (both regular and reverse), source vertex $x$, target vertex $y$, time interval $t_{\alpha},t_{\omega}]$, a set of beer vertices $B$ and a list $T_{i}$ of time units for each of them.
2:The SBP distance from $x$ to $y$ within $t_{\alpha},t_{\omega}]$.
[…]
@@
8:   $a^{\prime}\leftarrow\min\{time\in T_{b}\mid time\geq a\}$ $\triangleright$ $\infty$ if it does not exist
@@
12:     Remove $(d,a)$ from $L_{b}$ $\triangleright$ impossible to get beer
@@
14:   Find $(d^{\prime},s)$ in $I_{b}$ with minimum $s$ such that $s\geq a$ $\triangleright$ find best fitting path